Suppose a stock's price is $38. and the comtinuously compounded interest rate is 6.5%. The stock does not pay dividends. a 3-month $40-strike European call costs $1.93, and a 3-month $40-strike put costs $3.84. In this situation, and arbitrager would....

(BUY or SELL) the stock
(BUY or SELL) the call
(BUY or SELL) the put
Answer #1

As per Put Call Parity, the prices of options with same strike price & expiry date are as follows:

Price of Call + PV of Exercise Price = Spot Price (Current Stock Price) + Price of Put

Therefore,

1.93 + [40*(e^(-0.065*3/12))] = 38 + P

1.93 + [40*(e^-0.01625)] = 38 + P

1.93 + [40*0.98385(from table)] = 38 + P

1.93 + 39.354 – 38 = P

Therefore, Price of Put = P = 3.284

Actual Price of Put > Theoretical Price. Therefore, Put is Overvalued.

Arbitrage Strategy:

As Put is Overvalued, Buy Call Option, Sell Stock Now, Sell Put Option